# General geometry realized by four-scalar model and application to $f(Q)$ gravity

### v1

@@ II Gravity coupled to two scalars
Here $S_{\mathrm{gravity}}$ is the action of the arbitrary gravity theory, $S_{\mathrm{matter}}$ is the action of matter, $A(\phi,\chi)$, $B(\phi,\chi)$, and $C(\phi,\chi)$ are functions of the two scalar fields $\phi$ and $\chi$, and $V(\phi,\chi)$ is the potential for these scalar fields $\phi$ and $\chi$ fields. In (II), the matter Lagrangian is expressed by $\mathcal{L}_{\mathrm{matter}}$. The gravitational coupling constant $\kappa$ is defined by using Newton’s gravitational constant $G$, $\kappa^{2}=8\pi G$, although we mainly use geometrized units, where the light speed $c$ and Newton’s gravitational constant $G$ are unity, $c=G=1$ throughout the paper.
By the variation of the action (II) with respect to the metric $g_{\mu\nu}$, we obtain the following Einstein equation,
@@
$\,+\frac{1}{2}\left\{A(\phi,\chi)\partial_{\mu}\phi\partial_{\nu}\phi+B(\phi,\chi)\left(\partial_{\mu}\phi\partial_{\nu}\chi+\partial_{\nu}\phi\partial_{\mu}\chi\right)+C(\phi,\chi)\partial_{\mu}\chi\partial_{\nu}\chi\right\}+\frac{1}{2}T_{\mathrm{matter}\,\mu\nu}\,.$ (2)
Here $T_{\mathrm{matter}\,\mu\nu}$ is the energy-momentum tensor of matter and $\mathcal{G}_{\mu\nu}$ is defined by the variation of the action $S_{\mathrm{gravity}}$ of the gravity sector as follows,
$\mathcal{G}^{\mu\nu}\equiv\frac{1}{\sqrt{-g}}\frac{\delta S_{\mathrm{gravity}}}{\delta g_{\mu\nu}}\,.$ (3)
@@
As discussed in $A$ and/or $C$ become often negative, and therefore $\phi$ and/or $\chi$ become ghosts. The elimination of the ghosts can be done by imposing constraints by introducing the Lagrange multiplier fields $\lambda_{\phi}$ and $\lambda_{\chi}$. We modify the action $S_{\phi\chi}$ in (II) by $S\to S+S_{\lambda}$, where
$S_{\lambda}=\int d^{4}x\sqrt{-g}\left[\lambda_{\phi}\left(\mathrm{e}^{-2\nu(t=\phi,r=\chi)}\partial_{\mu}\phi\partial^{\mu}\phi+1\right)+\lambda_{\chi}\left(\mathrm{e}^{-2\lambda(t=\phi,r=\chi)}\partial_{\mu}\chi\partial^{\mu}\chi-1\right)\right]\,.$ (7)
@@
$0=\mathrm{e}^{-2\nu(t=\phi,r=\chi)}\partial_{\mu}\phi\partial^{\mu}\phi+1\,,\quad 0=\mathrm{e}^{-2\lambda(t=\phi,r=\chi)}\partial_{\mu}\chi\partial^{\mu}\chi-1\,.$ (8)
@@
$\partial_{t}\left(\mathrm{e}^{-2\nu(t,r)}\delta\phi\right)=\partial_{r}\left(\mathrm{e}^{-2\lambda(t,r)}\delta\chi\right)=0\,.$ (10)
Eq. (22) tells that by imposing the initial condition $\delta\phi=0$ and by imposing the boundary condition $\delta\chi\to 0$ when $r\to\infty$, we find that both of $\delta\phi$ and $\delta\chi$ vanish in the whole spacetime, $\delta\phi=0$ and $\delta\chi=0$. This tells that both $\phi$ and $\chi$ are non-dynamical or frozen degrees of freedom.
@@
$2\mathcal{G}_{\theta\theta}=2\mathcal{G}_{\phi\phi}\sin^{2}\theta=$ $\,r^{2}\left(\frac{A}{2}\mathrm{e}^{-2\nu}-\frac{C}{2}\mathrm{e}^{-2\lambda}-V\right)+r^{2}p\,,$
@@
$0=$ $\,\mathcal{G}_{t\theta}=\mathcal{G}_{\theta t}=\mathcal{G}_{r\theta}=\mathcal{G}_{\theta r}=\mathcal{G}_{t\phi}=\mathcal{G}_{\phi t}=\mathcal{G}_{r\phi}=\mathcal{G}_{\phi r}\,.$ (11)
Here we denote differentiations with respect to the time coordinate $t$ and radial coordinate $r$ by the dot “$\dot{\ }$” and prime “${}^{\prime}$”, respectively. We also assume that the matter is a perfect fluid and $\rho$ and $p$ are the energy density and the pressure of matter, defined by,
$T_{\mathrm{matter}\,tt}=\rho=-g_{tt}\rho\ ,\quad T_{\mathrm{matter}\,ij}=pg_{ij}\,.$ (12)
In the case of Einstein’s gravity, we find that $\mathcal{G}_{t\theta}=\mathcal{G}_{\theta t}$, $\mathcal{G}_{r\theta}=\mathcal{G}_{\theta r}$, $\mathcal{G}_{t\phi}=\mathcal{G}_{\phi t}$, and $\mathcal{G}_{r\phi}=\mathcal{G}_{\phi r}$ vanish trivially. As we will see later, however, in the case of $f(Q)$ gravity $\mathcal{G}_{r\theta}=\mathcal{G}_{\theta r}$ does not vanish and the equation $\mathcal{G}_{r\theta}=\mathcal{G}_{\theta r}=0$ gives non-trivial constraints on the model.
When $\mathcal{G}_{t\theta}=\mathcal{G}_{\theta t}$, $\mathcal{G}_{r\theta}=\mathcal{G}_{\theta r}$, $\mathcal{G}_{t\phi}=\mathcal{G}_{\phi t}$, and $\mathcal{G}_{r\phi}=\mathcal{G}_{\phi r}$ vanish trivially, the equations (II) can be algebraically solved with respect to $A$, $B$, $C$. and $V$ as follows,
$A=2\left(\mathcal{G}_{tt}+\frac{\mathrm{e}^{2\nu}}{r^{2}}\mathcal{G}_{\theta\theta}\right)-\mathrm{e}^{2\nu}\left(\rho+p\right)\,,\quad B=2\mathcal{G}_{tr}\,,\quad C=2\left(\mathcal{G}_{rr}-\frac{\mathrm{e}{2\lambda}}{r^{2}}\mathcal{G}_{\theta\theta}\right)\,,$
@@ III Four-scalar model
Because in the gravity theory where any of $\mathcal{G}_{t\theta}=\mathcal{G}_{\theta t}$, $\mathcal{G}_{r\theta}=\mathcal{G}_{\theta r}$, $\mathcal{G}_{t\phi}=\mathcal{G}_{\phi t}$, and $\mathcal{G}_{r\phi}=\mathcal{G}_{\phi r}$ does not vanish, even if we use the two-scalar model in Section II, we cannot construct a model which realizes a given spherically symmetric spacetime. In order to solve the problem, we now consider further extension of the model by using four scalar fields. We also show that we can realize any geometry by using the model.
@@
Here the kinetic coefficients $A_{\left(\rho\sigma\right)}\left(\phi^{\left(\tau\right)}\right)$ and the potential $V\left(\phi^{\left(\rho\right)}\right)$ are functions of the scalar fields $\phi^{\left(\rho\right)}$. In $S_{\lambda}$, $\lambda^{(\rho)}$’s are Lagrange multiplier fields which give constraints as in (8),
@@
For the perturbation $\delta\phi^{\left(\rho\right)}$, the constraints in (15) give
$0=\frac{\partial_{\xi}g^{\rho\rho}\left(x^{\sigma}\right)}{g^{\rho\rho}}\left(x^{\sigma}\right)\delta\phi^{\left(\xi\right)}+\partial_{\xi}\delta\phi^{\left(\xi\right)}\,.$ (22)
We should note that we have not summed the equations with respect to $\rho$. If there is a $\rho\neq\xi$ so that $\partial_{\xi}g^{\rho\rho}\left(x^{\sigma}\right)$ does not vanish, we obtain $\delta\phi^{\left(\xi\right)}=0$. Even if $\partial_{\xi}g^{\rho\rho}\left(x^{\sigma}\right)$ vanishes for all $\rho\neq\xi$, when $\rho=\xi$, we find $\partial_{\xi}\left(g^{\xi\xi}\delta\phi^{\left(\xi\right)}\right)=0$, that is, $g^{\xi\xi}\delta\phi^{\left(\xi\right)}$ does not depend on $x^{\xi}$. Then if $x^{\xi}$ is a space-like coordinate, if we impose $\delta\phi^{\left(\xi\right)}=0$ when $\left|x^{\xi}\right|\to\infty$, we find $\delta\phi^{\left(\xi\right)}=0$ everywhere. If $x^{\xi}$ is a time-like coordinate, if we impose $\delta\phi^{\left(\xi\right)}=0$ as an initial condition, we always find $\delta\phi^{\left(\xi\right)}=0$. Therefore $\delta\phi^{\left(\xi\right)}$ does not propagate and the ghosts do not appear.
We should note that it is straightforward to extend the above model into the model in general $D$ dimensions by using $D$ scalar fields.
@@ IV.1 Brief review
When both the torsion and the non-metricity equal zero, ${T^{\sigma}}_{\mu\nu}={Q^{\sigma}}_{\mu\nu}=0$, the connection ${\Gamma^{\sigma}}_{\mu\nu}$ reduces to the Levi-Civita connection ${{\tilde{\Gamma}}^{\sigma}}_{\mu\nu}$. In the Symmetric Teleparallel Equivalent of General Relativity (STEGR), the torsion ${T^{\sigma}}_{\mu\nu}$ and the curvature $R^{\lambda}_{\ \mu\rho\nu}$ given by the general affine connection ${\Gamma^{\sigma}}_{\mu\nu}$ in (23) vanish $R^{\lambda}_{\ \mu\rho\nu}\equiv{\Gamma^{\lambda}}_{\mu\nu,\rho}-{\Gamma^{\lambda}}_{\mu\rho,\nu}+{\Gamma^{\eta}}_{\mu\nu}\Gamma^{\lambda}_{\rho\eta}-{\Gamma^{\eta}}_{\mu\rho}\Gamma^{\lambda}_{\nu\eta}=0$. Although the STEGR gravity depends on both metric and connection satisfying the above constraints ${T^{\sigma}}_{\mu\nu}=R^{\lambda}_{\ \mu\rho\nu}=0$, the coincident gauge where ${\Gamma^{\sigma}}_{\mu\nu}=0$ are often used although the bases of the gauge are not so clear but the gauge condition could correspond to the choice of the coordinate system. In the coincident gauge, the non-metricity tensor in Eq. (28) reduces to the following form,
@@
$\mathcal{G}_{\mu\nu}=$ $\,\frac{1}{\sqrt{-g}}g_{\mu\rho}g_{\nu\sigma}\frac{\delta S}{\delta g_{\rho\sigma}}$

### v2

@@ I Introduction
As conventions, the Greek indices $\mu,\nu=0,1,2,3$ and the Lattin indices $i,j=1,2,3$. We also choose the signature of the metric as (,+,+,+)(-,+,+,+)( - , + , + , + ).
@@ II Gravity coupled to two scalars
Here $S_{\mathrm{gravity}}$ is the action of the arbitrary gravity theory, $S_{\mathrm{matter}}$ is the action of matter, $A(\phi,\chi)$, $B(\phi,\chi)$, and $C(\phi,\chi)$ are functions of $\phi$ and $\chi$, and $V(\phi,\chi)$ is the potential for $\phi$ and $\chi$.
By the variation of the action (II) with respect to the metric $g_{\mu\nu}$, we obtain the following equation corresponding to the Einstein equation,
@@
$\,+\frac{1}{2}\left\{A(\phi,\chi)\partial_{\mu}\phi\partial_{\nu}\phi+B(\phi,\chi)\left(\partial_{\mu}\phi\partial_{\nu}\chi+\partial_{\nu}\phi\partial_{\mu}\chi\right)+C(\phi,\chi)\partial_{\mu}\chi\partial_{\nu}\chi\right\}+\frac{1}{2}\left(T_{\mathrm{matter}}\right)_{\mu\nu}\,.$ (2)
Here $\left(T_{\mathrm{matter}}\right)_{\mu\nu}$ is the energy-momentum tensor of matter defined by $\left(T_{\mathrm{matter}}\right)^{\mu\nu}\equiv\frac{2}{\sqrt{-g}}\frac{\delta S_{\mathrm{matter}}}{\delta g_{\mu\nu}}$ and $\mathcal{G}_{\mu\nu}$ is defined by the variation of the action $S_{\mathrm{gravity}}$ of the gravity sector as follows,
$\mathcal{G}^{\mu\nu}\equiv-\frac{1}{\sqrt{-g}}\frac{\delta S_{\mathrm{gravity}}}{\delta g_{\mu\nu}}\,.$ (3)
@@
As discussed in Nojiri:2020blr ; Nojiri:2023dvf ; Nojiri:2023zlp ; Elizalde:2023rds ; Nojiri:2023ztz , $A$ and/or $C$ become often negative, and therefore $\phi$ and/or $\chi$ become ghosts. The elimination of the ghosts can be done by imposing constraints by introducing the Lagrange multiplier fields $\lambda_{\phi}$ and $\lambda_{\chi}$. We modify the action $S_{\phi\chi}$ in (II) by $S\to S+S_{\lambda}$, where
$S_{\lambda}=\int d^{4}x\sqrt{-g}\left[\lambda_{\phi}\left(\mathrm{e}^{2\nu(t=\phi,r=\chi)}\partial_{\mu}\phi\partial^{\mu}\phi+1\right)+\lambda_{\chi}\left(\mathrm{e}^{2\lambda(t=\phi,r=\chi)}\partial_{\mu}\chi\partial^{\mu}\chi-1\right)\right]\,.$ (7)
@@
$0=\mathrm{e}^{2\nu(t=\phi,r=\chi)}\partial_{\mu}\phi\partial^{\mu}\phi+1\,,\quad 0=\mathrm{e}^{2\lambda(t=\phi,r=\chi)}\partial_{\mu}\chi\partial^{\mu}\chi-1\,.$ (8)
@@
$\partial_{t}\left(\mathrm{e}^{2\nu(t=\phi,r=\chi)}\delta\phi\right)=\partial_{r}\left(\mathrm{e}^{2\lambda(t=\phi,r=\chi)}\delta\chi\right)=0\,.$ (10)
Eq. (10) tells that by imposing the initial condition $\delta\phi=0$ (because $\phi$ corresponds to the time coordinate) and by imposing the boundary condition $\delta\chi\to 0$ when $r\to\infty$ (because $\chi$ corresponds to the radial coordinate), we find that both of $\delta\phi$ and $\delta\chi$ vanish in the whole spacetime, $\delta\phi=0$ and $\delta\chi=0$ not only at the initial surface for $\delta\phi$ and boundary surface for $\delta\chi$. This tells that both $\phi$ and $\chi$ are non-dynamical or frozen degrees of freedom.
@@
$2\mathcal{G}_{\vartheta\vartheta}=\frac{2\mathcal{G}_{\varphi\varphi}}{\sin^{2}\vartheta}=$ $\,r^{2}\left(\frac{A}{2}\mathrm{e}^{-2\nu}-\frac{C}{2}\mathrm{e}^{-2\lambda}-V\right)+r^{2}p\,,$
@@
$0=$ $\,\mathcal{G}_{t\vartheta}=\mathcal{G}_{\vartheta t}=\mathcal{G}_{r\vartheta}=\mathcal{G}_{\vartheta r}=\mathcal{G}_{t\varphi}=\mathcal{G}_{\varphi t}=\mathcal{G}_{r\varphi}=\mathcal{G}_{\varphi r}\,.$ (11)
Here we assume that the matter is a perfect fluid and $\rho$ and $p$ are the energy density and the pressure of matter, defined by,
$\left(T_{\mathrm{matter}}\right)_{tt}=\rho=-g_{tt}\rho\ ,\quad\left(T_{\mathrm{matter}}\right)_{ij}=pg_{ij}\,.$ (12)
In the case of Einstein’s gravity, we find that $\mathcal{G}_{t\vartheta}=\mathcal{G}_{\vartheta t}$, $\mathcal{G}_{r\vartheta}=\mathcal{G}_{\vartheta r}$, $\mathcal{G}_{t\varphi}=\mathcal{G}_{\varphi t}$, and $\mathcal{G}_{r\varphi}=\mathcal{G}_{\varphi r}$ vanish trivially. As we will see later, however, in the case of $f(Q)$ gravity $\mathcal{G}_{r\vartheta}=\mathcal{G}_{\vartheta r}$ does not vanish and the equation $\mathcal{G}_{r\vartheta}=\mathcal{G}_{\vartheta r}=0$ gives non-trivial constraints on the model.
When $\mathcal{G}_{t\vartheta}=\mathcal{G}_{\vartheta t}$, $\mathcal{G}_{r\vartheta}=\mathcal{G}_{\vartheta r}$, $\mathcal{G}_{t\varphi}=\mathcal{G}_{\varphi t}$, and $\mathcal{G}_{r\varphi}=\mathcal{G}_{\varphi r}$ vanish trivially, the equations (II) can be algebraically solved with respect to $A$, $B$, $C$. and $V$ as follows,
$A=2\left(\mathcal{G}_{tt}+\frac{\mathrm{e}^{2\nu}}{r^{2}}\mathcal{G}_{\vartheta\vartheta}\right)-\mathrm{e}^{2\nu}\left(\rho+p\right)\,,\quad B=2\mathcal{G}_{tr}\,,\quad C=2\left(\mathcal{G}_{rr}-\frac{\mathrm{e}^{2\lambda}}{r^{2}}\mathcal{G}_{\vartheta\vartheta}\right)\,,$
@@ III Four-scalar model
Because in the gravity theory where any of $\mathcal{G}_{t\vartheta}=\mathcal{G}_{\vartheta t}$, $\mathcal{G}_{r\vartheta}=\mathcal{G}_{\vartheta r}$, $\mathcal{G}_{t\varphi}=\mathcal{G}_{\varphi t}$, and $\mathcal{G}_{r\varphi}=\mathcal{G}_{\varphi r}$ does not vanish, even if we use the two-scalar model in Section II, we cannot construct a model which realizes a given spherically symmetric spacetime. In order to solve the problem, we now consider further extension of the model by using four scalar fields. We also show that we can realize any geometry by using the model.
@@
Here the kinetic coefficients $A_{\left(\rho\sigma\right)}\left(\phi^{\left(\tau\right)}\right)$ and the potential $V\left(\phi^{\left(\rho\right)}\right)$ are functions of the scalar fields $\phi^{\left(\rho\right)}$. We should note that the indices $\rho\sigma$ in $A_{(\rho\sigma)}$ etc. do not represent the indices of the tensor. In order to distinguish them, we use the notation with parentheses. We used Greek indices, however, because we will identify $\phi^{(\rho)}$ with $x^{\rho}$ later on. For simplicity, we exclude the contributions from matter since their inclusion, while straightforward, leads to more complex expressions. The factor $\frac{1}{g^{\rho\rho}\left(x^{\sigma}=\phi^{\left(\sigma\right)}\right)}$ in $S_{\lambda}$ might look to break the diffeomorphism invariance. We should note, however, that this factor is not a component of the metric but a function of the scalar fields and therefore this factor is consistent with the diffeomorphism invariance.
In $S_{\lambda}$, $\lambda^{(\rho)}$’s are Lagrange multiplier fields which give constraints as in (8),
@@
For the perturbation $\delta\phi^{\xi}$, the constraints in Eq. (15) give
$0=2g^{\xi\nu}\left(x^{\sigma}=\phi^{(\sigma)}\right)\partial_{\nu}\delta\phi^{(\xi)}-\sum_{\zeta}\delta\phi^{(\zeta)}\partial_{\zeta}g^{\xi\xi}\left(x^{\sigma}=\phi^{(\sigma)}\right)\,.$ (22)
Here, we have not summed the equations with respect to $\xi$. In Eq. (22), as long as we consider the perturbation from the background, we may replace $\phi^{(\sigma)}$ inside $g^{\xi\nu}\left(x^{\sigma}=\phi^{(\sigma)}\right)$ and $g^{\xi\xi}\left(x^{\sigma}=\phi^{(\sigma)}\right)$ with $x^{\sigma}$. For a space-like coordinate $x^{\xi}$, if we impose $\delta\phi^{(\xi)}=0$ when $\left|x^{\xi}\right|\to\infty$, and for a time-like coordinate $x^{\xi}$, if we impose $\delta\phi^{(\xi)}=0$ as an initial condition, we always find $\delta\phi^{(\xi)}=0$. Therefore, $\delta\phi^{(\xi)}$ does not propagate, and thus the ghosts do not appear.
In the above discussion, when we consider the perturbation in the fixed background, the scalar fields $\phi^{\left(\xi\right)}$ are not dynamical. In the perturbation in (21), we do not include the perturbation of the metric. In general, some components in the metric may mix with the scalar fields $\phi^{\left(\xi\right)}$ because the scalar fields couple with gravity, of course. In fact, Eq. (22) is modified as $0=2g^{\xi\nu}\left(x^{\sigma}\right)\partial_{\nu}\delta\phi^{(\xi)}-\sum_{\zeta}\delta\phi^{(\zeta)}\partial_{\zeta}g^{\xi\xi}\left(x^{\sigma}\right)-\delta g^{\xi\xi}\left(x^{\sigma}\right)$. The perturbation $\delta g^{\xi\xi}$ includes the scalar mode and the tensor mode with $+$-polarization. In Einstein’s gravity or even $f(Q)$ gravity, the scalar modes are not dynamical nor do not propagate. As a dynamical system, because the scalar mode in the metric is not dynamical, this mode does not have conjugate momenta. Although there appear conjugate momenta for the scalar fields $\phi^{(\xi)}$’s, they do not propagate by the constraints. Therefore even if there is a mixture of the scalar modes, the dynamical modes could not appear. Furthermore, although the scalar fields $\phi^{(\xi)}$’s couple with $+$-tensor mode, the scalar perturbation $\delta\phi^{(\xi)}$’s could not mix with the tensor mode. At least, it could not generate a new propagating mode although we need more detailed analyses, which could be checked in future work including more general gravity theory.
As in this paper, some extensions of the mimetic gravity have been proposed and investigated Takahashi:2017pje ; Langlois:2018jdg ; Ganz:2018mqi and some instabilities have been also found. This may suggest that we need more detailed analyses in order to check the consistencies of the model, which could be also considered in the future works.
Finally, we should note that it is straightforward to extend the above model into the model in general $D$ dimensions by using $D$ scalar fields.
@@ IV.1 Brief review
When both the torsion and the non-metricity equal zero, ${T^{\sigma}}_{\mu\nu}={Q^{\sigma}}_{\mu\nu}=0$, the connection ${\Gamma^{\sigma}}_{\mu\nu}$ reduces to the Levi-Civita connection ${{\tilde{\Gamma}}^{\sigma}}_{\mu\nu}$. In the Symmetric Teleparallel Equivalent of General Relativity (STEGR), the torsion ${T^{\sigma}}_{\mu\nu}$ and the curvature $R^{\lambda}_{\ \mu\rho\nu}$ given by the general affine connection ${\Gamma^{\sigma}}_{\mu\nu}$ in (23) vanish,
$R^{\lambda}_{\ \mu\rho\nu}\equiv{\Gamma^{\lambda}}_{\mu\nu,\rho}-{\Gamma^{\lambda}}_{\mu\rho,\nu}+{\Gamma^{\eta}}_{\mu\nu}{\Gamma^{\lambda}}_{\rho\eta}-{\Gamma^{\eta}}_{\mu\rho}{\Gamma^{\lambda}}_{\nu\eta}=0\,.$ (30)
A solution of Eq. (30) is given by using four fields $\xi^{a}$ $\left(a=0,1,2,3\right)$ as follows,
${\Gamma^{\rho}}_{\mu\nu}=\frac{\partial x^{\rho}}{\partial\xi^{a}}\partial_{\mu}\partial_{\nu}\xi^{a}\,.$ (31)
As we mentioned, there is a problem with the number of degrees of freedom in $f(Q)$ gravity Hu:2022anq ; DAmbrosio:2023asf ; Hu:2023gui . Although the functional degrees of freedom in the connection are restricted by the curvature-free and torsion-free conditions, it is not so clear which could be valid in the equations given by the variation with respect to the connection. In order to avoid this problem, we regard the metric $g_{\mu\nu}$ and $\xi^{a}$ as independent fields as in Nojiri:2024zab .
Although the STEGR gravity depends on both metric and connection satisfying the above constraints ${T^{\sigma}}_{\mu\nu}=R^{\lambda}_{\ \mu\rho\nu}=0$, the coincident gauge where ${\Gamma^{\sigma}}_{\mu\nu}=0$ are often used although the bases of the gauge are not so clear but the gauge condition could correspond to the choice of the coordinate system. To realize the coincident gauge, we may use the gauge fixing action $S_{\mathrm{gf}}$ by introducing the auxiliary fields $\zeta_{a}$ as follows,
$S_{\mathrm{gf}}=\int d^{4}x\sqrt{-g}\zeta_{a}\left(\xi^{a}-x^{a}\right)\,,$ (32)
which violates the general covariance because the coincident gauge corresponds to the coordinate choice. We should note that $\zeta_{a}$ is nothing but the Nakanishi-Lautrup field Nakanishi:1966zz ; Lautrup:1967zz .
In the coincident gauge, the non-metricity tensor in Eq. (28) reduces to the following form,
@@
$\mathcal{G}_{\mu\nu}=$ $\,\frac{1}{\sqrt{-g}}g_{\mu\rho}g_{\nu\sigma}\frac{\delta S_{f(Q)}}{\delta g_{\rho\sigma}}$
@@
We should note that although we chose the coincident gauge, there are equations given by the variation of the total action with respect to the connection. In our formulation, the connection is expressed by $\xi^{a}$. Therefore if the matter does not couple with $\xi^{a}$, we find
$0=\frac{1}{\sqrt{-g}}\frac{\delta\left(S_{f(Q)}+S_{\mathrm{gf}}\right)}{\delta\xi^{a}}=X_{a}+\zeta_{a}\,,\quad X_{a}\equiv\frac{1}{\sqrt{-g}}\frac{\delta S_{f(Q)}}{\delta\xi^{a}}\,,$ (37)
which can be always solved as $\zeta_{a}=-X_{a}$. We should note $X_{a}$ does not vanish in the general spacetime as in the spherically symmetric spacetime or the Friedmann-Lemaître-Robertson-Walker (FLRW) spacetime. Therefore if there is not the gauge fixing action $S_{\mathrm{gf}}$, Eq. (37) is not satisfied $X_{a}\neq 0$. We should note that $X_{a}$ corresponds to $\nabla_{\mu}\nabla_{\nu}\left(\sqrt{-g}f^{\prime}(Q){P^{\mu\nu}}_{\lambda}\right)$. In the previous research, the equation $\nabla_{\mu}\nabla_{\nu}\left(\sqrt{-g}f^{\prime}(Q){P^{\mu\nu}}_{\alpha}\right)=0$ was often imposed and it generates some conflictions between the coincident gauge and the spacetime. These conflictions occur because the gauge fixing action $S_{\mathrm{gf}}$ is not included in the total action. We should also note that $\frac{1}{\sqrt{-g}}g_{\mu\rho}g_{\nu\sigma}\frac{\delta S_{\mathrm{gf}}}{\delta g_{\rho\sigma}}=0$ because $\xi^{a}=x^{a}$, and therefore the gauge fixing action $S_{\mathrm{gf}}$ does not contribute to the equations given by the variation with respect to the metric.